OpenStudy (anonymous):

Solve (1/6x^2)=(1/2x)+(7/6x^2)

OpenStudy (loser66):

\[\huge\frac{1}{6x^2}=\frac{1}{2x}+\frac{7}{6x^2}\] \[~now,~ minus~ both~ sides ~by ~\frac{1}{6x^2}\]

OpenStudy (loser66):

what do you have?